OpenStudy (anonymous):

x/x+1 = x/x+4 Solve the proportion . What is the value(s) of x

OpenStudy (phi):

Did you try cross multiplying as your first step?

OpenStudy (anonymous):

x/x+1 = x/x+4 x*x+4 = x*x+1 x^2 +4x = x^2 + x subtract both sides and you get ? = 0

OpenStudy (anonymous):

\[\frac{x}{x+1}=\frac{x}{x+4}\]???

OpenStudy (anonymous):

Is that the question?

OpenStudy (anonymous):

\[x(x+1)=x(x+4)\]

OpenStudy (phi):

@patmyer1 what is wrong with x=0?